# : Adaptive Image Focus Optimization for Black-box MLLM

### v1

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have broadened the scope of vision-language tasks, excelling in applications like image captioning and interactive question-answering. However, these models struggle with accurately processing visual data, particularly in tasks requiring precise object recognition and fine visual details. Stringent token limits often result in the omission of critical information, hampering performance. To address these limitations, we introduce , a novel visual prompting mechanism designed to enhance MLLM performance while preserving essential visual details within token limits. features three key innovations: a prompt-aware strategy that dynamically highlights relevant image regions, a spatial-preserving orchestration schema that maintains object integrity, and a budget-aware prompting method that balances global context with crucial visual details. Comprehensive evaluations across multiple datasets demonstrate that consistently outperforms baseline methods, achieving up to a improvement in accuracy while significantly reducing token consumption.
@@ 1 Introduction
Vision-language understanding has witnessed remarkable progress with multimodal large language models (MLLMs) such as GPT-4o, Gemini Pro, and Claude 3.5Β [13, 8]. While these black-box MLLMs have demonstrated impressive capabilities in controlled settingsβeven being touted as having PhD-level skillsβour large-scale investigation reveals a critical hallucination problem: their inherent small object blindness and limited visibility lead to catastrophic failures in real-world visual reasoning tasks.
Through systematic analysis, we identify two fundamental limitations that create this problem. First, despite their sophisticated architectures, current black-box MLLMs process images with uniform attention, lacking the human-like ability to focus on relevant regions while maintaining spatial awareness. As shown in FigureΒ 1, this limitation leads to surprising failures in entry-level tasks like object countingβa problem that persists across different models and prompting strategies. Our investigation reveals that this is not merely an implementation issue but a fundamental limitation of how these models process visual information.
Second, we identify a critical tension between token budgets and visual fidelity in black-box MLLMs. While commodity black-box MLLM has treated token limitations, typically fixed token counts, as a simple engineering constraint to ensure computational efficiency and user fairnessΒ [4], our analysis reveals it as a fundamental resource allocation problem. As demonstrated in FigureΒ 2, the standard approach of uniform downsampling leads to catastrophic loss of fine detailsβa problem that cannot be solved by better downsampling algorithms or prompt engineering alone.
This investigation leads us to formulate a novel problem in visual prompting: How can we achieve region-aware visual processing within the strict token constraints of black-box MLLMs? This question challenges the conventional view that visual prompting is merely a simple input formatting task. Instead, we show that effective visual prompting requires solving three interconnected challenges: Region Selection: How to identify and prioritize task-relevant regions without access to model internals; Spatial Preservation: How to maintain structural relationships while maximizing detail preservation; Budget Optimization: How to allocate limited tokens between global context and local details.
Based on these insights, we propose , a visual prompting framework that addresses these challenges through three technical innovations:
@@
A prompt-aware visual emphasis mechanism that enables selective attention without modifying model architecture
@@
A spatial-preserving orchestration schema that maintains global context while preserving local details
@@
A budget-aware region selection strategy that optimally allocates tokens across image regions
In a comprehensive evaluation across datasets such as VstarΒ [28], CVBenchΒ [24], and RealworldQAΒ [29], consistently outperformed baseline methods. Notably, in the Vstar dataset, -Patches achieved a 26.9% accuracy improvement over the baseline, while in RealWorldQA, -Adaptive outperformed the baseline by 12.1%. Alongside these accuracy gains, significantly reduced token usage. For example, on the TerraIncognita dataset, achieved 6.4% higher accuracy with a 67% reduction in token consumption compared to the baseline. These results confirm that not only addresses the limitations of visual processing in black-box MLLMs but also enhances efficiency. Moreover, across APIs like GPT-4o111https://platform.openai.com/, Gemini-1.5Pro222https://gemini.google.com/, and Claude-3.5-Sonnet333https://anthropic.com/, demonstrated consistent improvements in both accuracy and token efficiency, solidifying its potential to optimize MLLM performance in real-world, high-resolution applications.
The contribution of this work can be summarized as follows: (1) We conduct a detailed investigation of GPT-4oβs image prompting strategy, exposing key limitations in its handling of visual inputs. (2) We introduce , a novel mechanism that addresses the challenge of preserving visual detail in black-box MLLMs while adhering to token constraints. (3) We present extensive experimental results across multiple datasets, demonstrating that achieves significant improvements in both accuracy and token efficiency, offering valuable insights into enhancing multimodal processing in constrained environments.
@@ 2 Pilot Experiments
One of the primary challenges faced by black-box MLLMs is their inability to process visual inputs efficiently, leading to diminished accuracy in fine-grained visual tasks. Models like GPT-4o often struggle with recognizing detailed or occluded objects, particularly when dealing with complex images. These limitations are compounded by token constraints, which restrict the amount of image data that can be processed in a single prompt. Furthermore, according to the Vision pricing calculator444https://openai.com/api/pricing/, GPT handle images by resizing and splitting them into basic units of 512Γ512 pixels. Each of these units corresponds to 170 tokens. This method of processing not only imposes a strict limit on the image resolution but also increases the computational overhead due to the additional tokens generated from splitting. As a result, vital visual details may be lost when images are downsampled or resized to fit within the token limits, leading to poor performance on tasks that require precise visual grounding.
@@
To evaluate this issue, we conducted a series of pilot experiments using GPT-4o-0513 on the Vstar-Bench dataset. This dataset challenges MLLMs to accurately identify detailed objects within high-resolution images, making it an ideal test for the modelβs capacity to handle fine-grained visual information. The experiments compared three different image processing strategies: (1) an unprocessed prompt (Unaltered Input), where the image is fed to the model in its original form; (2) a prompt where the image is cropped to focus on the target object (Image Crop); and (3) a prompt where the cropped image is further enlarged to emphasize the most relevant visual features (Zoomed Crop). Both the Image Crop and Zoomed Crop methods were constrained to fit within GPT-4oβs patch size limit of 512x512 pixels.
As shown in TableΒ 1, the Zoomed Crop method significantly outperformed the others, achieving an accuracy of 0.76 with a token usage of 270. In comparison, the Unaltered Input method, despite processing the entire image, only achieved an accuracy of 0.64 while consuming 955 tokens. Similarly, the Image Crop method, although reducing the token count to 270, did not yield any improvement in accuracy compared to the unprocessed input.
These results highlight a fundamental problem in current black-box MLLMs: they fail to efficiently manage the trade-off between image resolution and token constraints. In cases like those presented by the Vstar-Bench dataset, where fine-grained visual information is critical, processing unaltered high-resolution images leads to excessive token consumption without improving accuracy. While the Image Crop method reduces token usage, it fails to improve performance because simply cropping an image without emphasizing key details does not provide sufficient context for the model to interpret the visual input accurately.
The superior performance of the Zoomed Crop method underscores the importance of vision enhancement techniques in black-box MLLMs. By focusing on the most relevant portions of the image, Zoomed Crop preserves critical details while remaining within token limits, enabling the model to interpret detailed visual inputs more effectively. This approach resolves a common issue faced by black-box MLLMs, where downscaling or cropping images to meet token requirements often leads to a loss of essential information, reducing the overall effectiveness of the model.
Our experiments reveal that without adaptive techniques like Zoomed Crop, black-box MLLMs struggle to process high-resolution images efficiently, limiting their performance on tasks that require precise visual recognition. These findings demonstrate the necessity of vision enhancement strategies to address the inherent limitations of token-constrained MLLMs.
@@ 3.1 Multimodal LLMs: Open-Source and Black-Box Models
The integration of visual and textual modalities in large language models (LLMs) has led to significant advancements in multimodal models (MLLMs) like GPT-4o, Gemini Pro and Claude3-Sonnet. These models rely on effective visual encoding strategies to bridge the gap between language and vision. Approaches such as CLIPΒ [31] align visual and language embeddings through contrastive learning, while models like FlamingoΒ [1] and BLIP-2Β [6] use cross-attention mechanisms or pretraining modules to link vision encoders with LLMs. However, these methods often rely on fixed low-resolution inputs (e.g., 224x224), limiting their ability to process high-resolution images or non-standard aspect ratiosΒ [16], which hampers performance on fine-grained tasks such as OCR and small object detection
In contrast, open-source multimodal modelsΒ [15, 30, 36, 12, 37] allow for architectural modifications and fine-tuning to accommodate any-resolution inputs. However, black-box MLLMs such as GPT-4o and Gemini Pro, which impose strict token limits for computational efficiency, require alternative solutions. The need to downsample or crop images to meet these constraints often results in the loss of crucial visual details, particularly in tasks requiring detailed visual understanding. While position embedding interpolationΒ [2, 26, 19, 9, 5] and patch-based croppingΒ [30, 12] widely adpoted in open-soure models offer promising directions for any aspect ratio and any-resolution image processing, they are not applicable to black-box models, where architectural changes and extra training/fine-tuning are not permitted.
@@ 3.2 Object Detection
Traditional object detection models, such as Faster R-CNNΒ [22] and YOLOΒ [21], effectively identify and localize objects within predefined categories. However, they struggle with open-set scenarios, where novel objects not seen during training need to be detected.
Recent advances address this limitation through open-set detection models that leverage natural language processing. For instance, OV-DETRΒ [34] integrates CLIP with object detection to generate category-specific bounding boxes from textual prompts, enabling detection in open-world settings. Similarly, GLIPΒ [14] reframes detection as a grounding problem, improving alignment between visual regions and textual descriptions. DetCLIPΒ [32] extends this further using pseudo labels from large-scale captioning datasets, enhancing generalization. Grounding DINOΒ [17], built on the DETR frameworkΒ [3], also advances open-set detection through natural language integration.
In addition, SAMΒ [11] and SAM-2Β [20] offer zero-prompt or minimal-prompt segmentation for arbitrary objects but lack robust text-prompt handling. EVF-SAMΒ [35] overcomes this by extending SAMβs capabilities to better manage complex text-based object segmentation.
@@ 4 Method Overview
Inspired by the observation derived from our pilot experiments, we propose , a comprehensive visual prompting mechanism designed to effectively address the loss of detail in images that occurs during the naive resizing process in current black box multimodal LLMs, such as GPT-4 and Gemini 1.5. As illustrated in FigureΒ 3, our mechanism comprises three key components: (1) A prompt-aware visual emphasizer that allocates high-fidelity image slices based on prompt texts to facilitate efficient and focused visual encoding; (2) A spatial-preserving encoding schema that consolidates the collected image slices while maintaining their relative spatial positions to create a condensed visual input; (3) A budget-aware prompting strategy that maximizes the accuracy of results obtained from the black box models while fits the budget requirement from users.
@@ 4.1 Prompt-aware Visual Emphasizer
In addition to basic stopword removal, we utilize dependency parsingΒ [23, 7] to analyze the syntactic structure of the prompt. This deeper analysis identifies core entities and relationships, such as subject-object pairs and action verbs, which are crucial for interpreting the userβs intent. By focusing on these core semantic elements, we ensure that the visual emphasis aligns precisely with the underlying meaning of the prompt.
@@
Multi-Scale Emphasizing Algorithm: Given a key object term extracted from the text prompt, the Multi-Scale Emphasizing AlgorithmΒ 1 utilizes a state-of-the-art object detection model to localize the corresponding object in the image prompt. In our experiments, we primarily employ GroundingDINOΒ [17] as our localization model.
The encoder in such models typically downsamples the input image to a resolution of or , potentially resulting in information loss when localizing the target object at a coarse granularity. To address this limitation, we propose a Multi-Scale Emphasizing Algorithm that processes the original image at multiple resolutions. The algorithm divides the input image into patches at various granularities, e.g. , , and beyond. For each generated patch, we apply the object detection model to localize the target object. The algorithm retains bounding boxes returned by the model that exceed a predefined confidence threshold. These high-confidence bounding boxes collectively form the output of our algorithm, providing a comprehensive multi-scale representation of the target objectβs location.

### v2

# : Adaptive Visual Token Allocation for Black-box MLLM

@@ Abstract
Multimodal large language models (MLLMs) such as GPT-4o, Gemini Pro, and Claude 3.5 have enabled unified reasoning over text and visual inputs, yet they often hallucinate in real-world scenariosβespecially when small objects or fine spatial context are involved. We pinpoint two core causes of this failure: the absence of region-adaptive attention and inflexible token budgets that force uniform downsampling, leading to critical information loss. To overcome these limitations, we introduce , a visual prompting framework that delivers token-efficient, detail-preserving image representations for black-box MLLMs. integrates (1) a prompt-aware emphasis module to highlight semantically relevant regions, (2) a spatial-preserving orchestration schema to maintain object relationships, and (3) a budget-aware strategy to adaptively allocate tokens between global context and local details. Extensive experiments on nine benchmarks and three commercial MLLMs demonstrate that boosts accuracy by up to 27% while cutting image token usage by up to 67%. Our approach establishes a principled methodology for robust, resource-aware multimodal understanding in settings where model internals are inaccessible.
@@ 1 Introduction
Vision-language understanding has emerged as a central challenge in multimodal artificial intelligence, with broad applications ranging from robotics and autonomous driving to scientific diagram analysis and human-computer interaction. Recent advances in multimodal large language models (MLLMs), such as GPT-4o, Gemini Pro, and Claude 3.5, have significantly pushed the boundaries of this field by enabling unified reasoning over text and visual inputsΒ li2024evcap; gu2024anomalygpt. These models have demonstrated impressive performance on a range of benchmarks and are increasingly perceived as possessing near-human or even PhD-level capabilities in controlled environments. However, despite these promising results, our large-scale empirical study uncovers a critical limitation that remains underexplored: MLLMs exhibit a pronounced hallucination tendency in real-world visual reasoning tasks, largely due to their inherent difficulty in perceiving small objects and limited field-of-view awareness. This βsmall object blindnessβ can lead to severe and often undetected reasoning failures, raising concerns about the deployment of MLLMs in real-world visual reasoning scenarios.
To investigate the root causes of hallucination in MLLMs, we conduct a systematic analysis and uncover two fundamental limitations in the visual processing pipeline of current black-box MLLMs. First, these models lack region-adaptive attention mechanisms, resulting in uniform spatial processing that disregards task-relevant visual saliency. Despite their architectural complexity, black-box MLLMs fail to emulate the human visual systemβs ability to focus selectively while preserving contextual structure. As illustrated in FigureΒ 1, this limitation leads to persistent failure even in elementary tasks such as object counting. Notably, the issue is consistent across different model families and prompting strategies, indicating a fundamental limitation rooted in the design of their visual tokenization and attention mechanisms, rather than implementation-specific flaws.
Second, we identify a critical tension between token budget constraints and visual fidelity. Black-box MLLMs are constrained by fixed token limits for both textual and visual inputsβan engineering choice originally motivated by inference efficiency and fairness considerationsΒ (chen2024spatialvlm). However, our analysis reveals that token budgeting in vision tasks is a core computational bottleneck that directly impacts perception granularity. As shown in FigureΒ 2, the commonly adopted strategy of uniform image downsampling leads to substantial loss of local detail, especially for small or spatially distributed objects. Crucially, we demonstrate that neither prompt engineering nor improved resizing algorithms can compensate for this representational loss.
These observations motivate us to formally define a new problem in visual prompting for MLLMs: How can region-aware, spatially coherent visual inputs be constructed under strict token constraints in black-box inference settings? We argue that solving this problem requires a paradigm shift from heuristic prompt design to principled visual token orchestration. We decompose this problem into three interdependent sub-tasks: (1) Region Selection β identifying and prioritizing task-relevant regions without model internals; (2) Spatial Preservation β retaining spatial relations and object continuity across selected regions; and (3) Budget-Aware Prompting β balancing the allocation of tokens between global context and fine-grained details.
To address these challenges, we present , a visual prompting framework designed for token-efficient, region-aware visual representation. comprises three integrated components:
@@
A prompt-aware visual emphasis module, which selectively highlights salient regions by manipulating image input structure, enabling focused attention without altering the model;
@@
A spatial-preserving orchestration scheme, which ensures that contextual coherence and relative object positioning are preserved across split or patched inputs;
@@
A budget-aware region selection strategy, which dynamically allocates token capacity based on content density and task relevance.
We evaluate on a suite of diverse benchmarks, including Vstar(vstar), CVBench(tong2024cambrian1), and RealWorldQAΒ (xAI2024), which collectively capture controlled, open-domain, and in-the-wild reasoning scenarios. consistently outperforms competitive baselines in both accuracy and efficiency. On the Vstar dataset, the -Patches variant achieves a 26.9% absolute accuracy gain, while on RealWorldQA, -Adaptive exceeds baseline performance by 12.1%. Moreover, in the TerraIncognita setting, our method reduces token usage by 67% while improving accuracy by 6.4%. These gains hold across multiple API providers, including GPT-4o111https://platform.openai.com/, Gemini-1.5Pro222https://gemini.google.com/, and Claude-3.5-Sonnet333https://anthropic.com/, demonstrating that generalizes well across model architectures and tokenization schemes.
In summary, this work makes the following contributions: (1) We provide the first in-depth empirical analysis of how token allocation and spatial attention impact visual reasoning performance in black-box MLLMs. (2) We formulate the token-constrained, region-aware visual prompting problem and identify its core design dimensions. (3) We propose , a general-purpose framework for token-efficient, spatially coherent visual prompting. (4) We demonstrate that improves both accuracy and efficiency across models, datasets, and domains, highlighting a new path toward robust, resource-aware multimodal understanding.
@@ 2 Pilot Experiments
A fundamental challenge in black-box multimodal large language models (MLLMs) is their inability to process high-resolution visual inputs efficiently under strict token budget constraints. In platforms such as GPT-4o, visual inputs are internally segmented into fixed-size image patches (typically 512Γ512 pixels), each corresponding to a predefined token cost (e.g., 170 tokens per patch)444https://openai.com/api/pricing/. This patch-based tokenization strategy, designed for balancing inference efficiency and fairness, introduces a critical trade-off: higher image resolution yields more detailed information but leads to rapid token consumption, limiting the number of patches a model can process. Conversely, reducing token usage through downsampling or cropping risks omitting essential fine-grained visual cues.
To quantify the impact of this trade-off, we conducted a series of pilot experiments using GPT-4o-0513 on the Vstar-Bench dataset, which requires precise visual grounding of small or occluded objects in complex scenes. We compare three visual prompting strategies: (1) Unaltered Input, where the original image is submitted directly without preprocessing; (2) Image Crop, Unless otherwise stated, all regions of interest (ROIs) used in this paper are obtained automatically at inference time using the multi-scale, prompt-aware open-vocabulary detector pipeline described in SectionΒ LABEL:sec:emphasizer. We do not use any ground-truth bounding boxes from the datasets when constructing ROIs, including the introductory examples in this section; and (3) Zoomed Crop, which magnifies the cropped region to maximize visual feature visibility while preserving the 512Γ512 input constraint. Each cropped ROI is normalized to a consistent visual scale before composition. Concretely, we resize each ROI such that its longer side equals 512 pixels while preserving aspect ratio. ROIs that are originally smaller than this scale are not up-scaled beyond their native resolution, in order to avoid excessive blurring or hallucinations. For large input images, we first crop ROIs using the detector and then apply this max-sideβ512 rule per ROI. This normalization ensures a consistent scale before spatial composition in SectionΒ 4.2.
The results in TableΒ 1 reveal several critical insights. While Unaltered Input provides full-scene context, it incurs excessive token cost (955 tokens), with limited performance gainsβsuggesting that high token usage alone does not guarantee better model understanding. In contrast, the Image Crop method reduces token usage substantially but fails to improve accuracy, likely due to the lack of visual emphasis and limited discriminative detail in the cropped region.
The Zoomed Crop strategy significantly outperforms both baselines, achieving 64% accuracy with only 270 tokens. This improvement highlights the importance of visual emphasis: by magnifying relevant regions, Zoomed Crop restores critical visual granularity lost in standard downsampling, while maintaining input compactness. Importantly, this approach respects the modelβs fixed patch budget without requiring architectural changes or access to internal weights.
These findings underscore a fundamental limitation in black-box MLLMs: existing visual prompting techniques fail to manage the resolutionβtoken trade-off effectively. NaΓ―ve approaches such as downscaling or region cropping without adaptive enhancement do not sufficiently preserve semantic or structural information. Our pilot experiments suggest that vision enhancement strategiesβparticularly those that incorporate selective magnificationβare essential for preserving task-relevant detail in constrained inference settings.
In subsequent sections, we build upon this insight to design , a visual prompting framework that generalizes the principles of adaptive emphasis, spatial preservation, and budget-aware token allocation across diverse visual reasoning tasks.
@@ 3.1 Multimodal LLMs: Open-Source and Black-Box Models
The integration of visual and textual modalities in large language models (LLMs) has led to significant advancements in multimodal models (MLLMs) like GPT-4o, Gemini Pro and Claude3-Sonnet. These models rely on effective visual encoding strategies to bridge the gap between language and vision. Approaches such as CLIPΒ (yangAttentiveMaskCLIP2023) align visual and language embeddings through contrastive learning, while models like FlamingoΒ (alayracFlamingoVisualLanguage2022) and BLIP-2Β (daiInstructBLIPGeneralpurposeVisionLanguage2023) use cross-attention mechanisms or pretraining modules to link vision encoders with LLMs. However, these methods often rely on fixed low-resolution inputs (e.g., 224x224), limiting their ability to process high-resolution images or non-standard aspect ratiosΒ (liuVisualInstructionTuning2023), which hampers performance on fine-grained tasks such as OCR and small object detection
In contrast, open-source multimodal modelsΒ (liLLaMAVIDImageWorth; xuLLaVAUHDLMMPerceiving2024; zhangLLaVAHDDivingHighResolution2024; liLLaVANeXTInterleaveTacklingMultiimage2024; zhaoMGLLaVAMultiGranularityVisual2024) allow for architectural modifications and fine-tuning to accommodate any-resolution inputs. However, black-box MLLMs such as GPT-4o and Gemini Pro, which impose strict token limits for computational efficiency, require alternative solutions. The need to downsample or crop images to meet these constraints often results in the loss of crucial visual details, particularly in tasks requiring detailed visual understanding. While position embedding interpolationΒ (baiQwenVLVersatileVisionLanguage2023; wangQwen2VLEnhancingVisionLanguage2024; luoFeastYourEyes2024; hongCogAgentVisualLanguage2023; chenMiniGPTv2LargeLanguage2023) and patch-based croppingΒ (xuLLaVAUHDLMMPerceiving2024; liLLaVANeXTInterleaveTacklingMultiimage2024) widely adpoted in open-soure models offer promising directions for any aspect ratio and any-resolution image processing, they are not applicable to black-box models, where architectural changes and extra training/fine-tuning are not permitted.
@@ 3.2 Object Detection
Traditional object detection models, such as Faster R-CNNΒ (renFasterRCNNRealTime2016) and YOLOΒ (redmonYouOnlyLook2016), effectively identify and localize objects within predefined categories. However, they struggle with open-set scenarios, where novel objects not seen during training need to be detected.
Recent advances address this limitation through open-set detection models that leverage natural language processing. For instance, OV-DETRΒ (zangOpenVocabularyDETRConditional2022) integrates CLIP with object detection to generate category-specific bounding boxes from textual prompts, enabling detection in open-world settings. Similarly, GLIPΒ (liGroundedLanguageImagePretraining2022) reframes detection as a grounding problem, improving alignment between visual regions and textual descriptions. DetCLIPΒ (yaoDetCLIPDictionaryEnrichedVisualConcept2022) extends this further using pseudo labels from large-scale captioning datasets, enhancing generalization. Grounding DINOΒ (liuGroundingDINOMarrying2023), built on the DETR frameworkΒ (carionEndEndObjectDetection2020), also advances open-set detection through natural language integration.
In addition, SAMΒ (kirillov2023segment) and SAM-2Β (ravi2024sam2) offer zero-prompt or minimal-prompt segmentation for arbitrary objects but lack robust text-prompt handling. EVF-SAMΒ (zhangEVFSAMEarlyVisionLanguage2021) overcomes this by extending SAMβs capabilities to better manage complex text-based object segmentation.
@@ 4 Method Overview
Building upon the findings of our pilot experiments, we introduce , a unified visual prompting framework that enables detail-preserving, token-efficient encoding of high-resolution visual inputs for black-box multimodal LLMs. Existing models such as GPT-4o and Gemini 1.5 often rely on uniform downsampling or tiling to fit image inputs into fixed-size patches, which leads to significant information loss and degraded performance on fine-grained vision tasks.
addresses this challenge through three tightly coupled modules, as shown in FigureΒ 3:
β’
Prompt-Aware Visual Emphasizer: extracts task-relevant image regions based on semantic cues derived from natural language prompts.
β’
Spatial-Preserving Orchestration Schema: reconstructs extracted regions into spatially faithful layouts that preserve object relationships and scene structure.
β’
Budget-Aware Prompting Strategy: manages token allocation under user-specified constraints by adapting the number and organization of image regions.
@@ 4.1 Prompt-aware Visual Emphasizer
In addition to basic stopword removal, we utilize dependency parsingΒ (parse1; parse2) to analyze the syntactic structure of the prompt. We use a lightweight dependency parser to identify noun phrases and their syntactic roles (e.g., question objects versus modifiers). Rather than constructing a full scene graph, we use these dependencies as an importance-ranking mechanism: (1) higher-priority noun phrases are passed to the open-vocabulary detector first, and (2) when multiple entities compete for limited detection budget, the ranking is used to break ties. In other words, dependency parsing is used for importance weighting and ordering prior to detection, not for explicit relation modeling. This deeper analysis identifies core entities and relationships, such as subject-object pairs and action verbs, which are crucial for interpreting the userβs intent. By focusing on these core semantic elements, we ensure that the visual emphasis aligns precisely with the underlying meaning of the prompt.
@@
Multi-Scale Emphasizing Algorithm: Given a key object term extracted from the text prompt, the Multi-Scale Emphasizing AlgorithmΒ 1 utilizes a state-of-the-art object detection model to localize the corresponding object in the image prompt. In our experiments, we primarily employ GroundingDINOΒ (liuGroundingDINOMarrying2023) as our localization model.
The encoder in such models typically downsamples the input image to a resolution of or , potentially resulting in information loss when localizing the target object at a coarse granularity. To address this limitation, we propose a Multi-Scale Emphasizing Algorithm that processes the original image at multiple resolutions. The algorithm divides the input image into patches at various granularities, e.g., , , and beyond. For each generated patch, we apply the object detection model to localize the target object. The algorithm retains bounding boxes returned by the model that exceed a predefined confidence threshold. These high-confidence bounding boxes collectively form the output of our algorithm, providing a comprehensive multi-scale representation of the target objectβs location.